multiply and simplify
(squareroot 6-2root3) (squareroot 6 +2root3)

Answer by fractalier(6550)

Just FOIL it out...we have
(sqrt(6) - 2*sqrt(3)) * (sqrt(6) + 2*sqrt(3)) =
sqrt(36) + 2*sqrt(18) - 2*sqrt(18) - 4*sqrt(9) =
6 - 12 =
-6